Whitman Company has just completed its first year of operations. The company’s absorption costing income statement for the year follows:

Whitman Company
Income Statement
Sales (39,000 units × $45.10 per unit) $ 1,758,900
Cost of goods sold (39,000 units × $22 per unit) 858,000
Gross margin 900,900
Selling and administrative expenses 448,500
Net operating income $ 452,400

The company’s selling and administrative expenses consist of $292,500 per year in fixed expenses and $4 per unit sold in variable expenses. The $22 unit product cost given above is computed as follows:

Direct materials $ 9
Direct labor 4
Variable manufacturing overhead 3
Fixed manufacturing overhead ($276,000 ÷ 46,000 units) 6
Absorption costing unit product cost $ 22

Required:

1. Redo the company’s income statement in the contribution format using variable costing.

2. Reconcile any difference between the net operating income on your variable costing income statement and the net operating income on the absorption costing income statement above.
Answer #1
1
Sales 1758900
Variable expenses:
Variable cost of goods sold 624000 =39000*(9+4+3)
Variable selling and administrative expenses 156000 =39000*4
780000
Contribution margin 978900
Fixed expenses:
Fixed manufacturing overhead 276000
Fixed selling and administrative expenses 292500
568500
Net operating income 410400
2
Variable costing net income(loss) 410400
Add: Fixed manufacturing overhead deferred in
inventory under absorption costing
42000 =7000*6
Absorption costing net operating income 452400